# TransientTrack: Advanced Multi-Object Tracking and Classification of Cancer Cells with Transient Fluorescent Signals

### v1

@@ 1 Introduction
To evaluate our approach we used a particularly challenging dataset of $309$ timelapse videos of human osteosarcoma-derived cancer cell-line (U-2 OS) in culture. Each video is $93$ hours long and the cells are tagged with three fluctuating signals, corresponding to the circadian rhythm, cell fitness, and cell cycle. In $117$ videos, cisplatin, a chemotherapeutic drug, was added to induce DNA damage. For a subset of these videos, we manually annotated cell positions, key cellular events (mitosis and apoptosis), and lineage relationships by linking each cell to its parent and daughter cells. After establishing the tracking performance of our method on a subset of annotated videos, we applied it to the full dataset to conduct a large-scale analysis revealing dose-dependent reductions in proliferation, delayed apoptosis onset, and characteristic lineage signatures reflecting how treatment disrupts cell-state inheritance. Together, these results demonstrate how our approach can uncover both global and lineage-specific dynamics in complex imaging datasets.
@@ 3.1 Data
Our dataset examines the proliferation status and cell cycle phase response of single cells to a chemotherapeutic drug cisplatin (granada2020effects). The dataset comprises $309$ videos, each containing $234$ frames at a size of $1024\times 1024$ pixels per frame. Frames were captured at $30$-minute intervals over a total duration of $117$ hours per video using a fluorescence timelapse microscopy setup. To monitor cellular dynamics, three distinct fluorescent markers — cell cycle reporter hGeminin-CFP, circadian clock reporter NR1D1::VNP, and cell fitness reporter p53-mKate — were employed, resulting in multi-channel images that capture transient, oscillatory signals over time, providing insight into the state and activity of cells. In $118$ of these videos, cisplatin was added after $48$ hours (frame $96$) in three dosage groups: high ($46$ videos, $13$ µM), medium ($36$ videos, $10$ µM), and low ($36$ videos, $7$ µM). Cisplatin affects DNA synthesis and DNA-damage repair, leading to cell division inhibition and ultimately cell death (yu2016cisplatin). We also have $36$ control videos where cisplatin was not added.
@@
We utilized $10$ of the annotated videos to evaluate tracking performance. The remaining annotated videos were used to train the detector. They were randomly split into training, validation, and test sets with a 0.80.8/0.10.1/$0.1$ ratio, respectively. For analysis, we utilized a subset of the whole dataset for which we have drug dosage information ($154$ videos), including all $234$ frames per video. We published the complete dataset at https://doi.org/10.5281/zenodo.17226123.
@@ 3.2.2 Matching
Our matching is carried out sequentially between two consecutive frames. During the first stage of matching, we match high-confidence detections of frame $t$ and $t+1$ with each other. We base our matching on the distance between the centroids of the detections and the similarity of their embeddings. For the distance, we use the Euclidean distance d(c,c^)d($\hat{c}$) which is the distance between the centroids $c_{\text{cent}}$ and $\hat{c}_{\text{cent}}$ of two cells $c$ and $\hat{c}$, respectively. To enforce embedding consistency during matching, we introduce a similarity threshold $\tau^{\text{sim}}$ applied to the cosine similarity between cell embeddings $c_{\text{emb}}$ and $\hat{c}_{\text{emb}}$:
@@ 4.2.1 CTC Evaluation
Regarding the CTC metrics (Table 1), TransientTrack achieved a DET score of $0.9393$, the LNK measure of $0.9154$ and a TRA score of $0.9362$. For reference, DET scores reported in the CTC challenge range from $0.876$ to $0.997$, LNK scores from $0.947$ to $1.000$, and TRA scores from $0.802$ to $0.994$ across different datasets and their best-performing methods. Although these values are not directly comparable due to substantial differences in the datasets and their difficulty, it is noteworthy that our method achieved scores within a similar range — except for the LNK score — even though it was evaluated on a considerably more challenging dataset.
Figure 2 presents the performance of TransientTrack across different drug dose groups, evaluated using the TRA metric and their respective number of ground truth tracks. The results indicate that our approach achieved consistently high performance across the high-, medium-, and low-dose groups, which contain an average of $62$, $83$, and $102$ tracks per video, respectively. However, performance slightly declines as cell density increases, particularly in the control group. This group includes the fewest videos, yet exhibits the highest average track count per video ($\sim$300). Despite this, TransientTrack maintains strong tracking accuracy, with average TRA scores of $0.9337$, $0.9386$, and $0.9441$ for the low-, medium-, and high-dose groups, respectively.
@@
CTC Metrics DET $0.9393\pm 0.0100$
TRA $0.9362\pm 0.0106$
LNK $0.9154\pm 0.0185$
MOT Metrics HOTA $0.7661\pm 0.0362$
DetA $0.7773\pm 0.0332$
AssA $0.7563\pm 0.0456$
MOTA $0.7955\pm 0.0424$
MOTP $0.9250\pm 0.0056$
IDF1 $0.8168\pm 0.0409$
@@ 4.2.2 MOT Evaluation
TransientTrack showed consistent MOT performance across all evaluated metrics (Table 1). For overall tracking accuracy, as measured by HOTA, it reached $0.766$, with detection accuracy (DetA) of $0.778$ and association accuracy (AssA) of $0.756$. The MOTA score of $0.796$ indicates that most cells were tracked correctly despite the challenging imaging conditions, although some errors from missed detections, false positives, or identity switches still occurred. In contrast, the MOTP score exceeded $0.925$, highlighting that once cells were detected, their positions were estimated with very high spatial precision. The IDF1 score of $0.817$ confirmed that object identities were maintained reliably in time, supporting the construction of accurate lineage trees and the detection of mitotic and apoptotic events. Taken together, these results suggest that TransientTrack excels at precise cell localization while occasionally facing difficulties in maintaining perfect temporal associations, an expected outcome in videos with transient fluorescent markers, heterogeneous signal quality, and frequent apoptotic events.
@@
TransientTrack $\mathbf{0.939\pm 0.0010}$ $\mathbf{0.915\pm 0.018}$ $\mathbf{0.936\pm 0.011}$
Only High-Conf. Detections $0.920\pm 0.015$ $0.880\pm 0.031$ $0.915\pm 0.017$
No Kalman Filter $0.934\pm 0.013$ $0.898\pm 0.027$ $0.930\pm 0.015$
None $0.908\pm 0.016$ $0.857\pm 0.033$ $0.902\pm 0.018$
@@ 4.3 Ablations
In order to assess the contribution of various features or our method to the accuracy, we conducted ablation experiments by selectively omitting them (Table 2). We considered four configurations: all features enabled, usage of only high-confidence detections, no track interpolation with Kalman Filter, and neither. In order to evaluate the different settings, we adhered to the CTC metrics only, since they are more suitable for capturing detection, tracking, and mitosis detection. Only using high-confidence detections led to a substantial drop in performance, with the mean scores for the metrics decreasing from $0.9393$ to $0.9203$ (DET), $0.9154$ to $0.8797$ (LNK), and from $0.9362$ to $0.9150$ (TRA). This underscores the importance of low-confidence detections as a fallback mechanism for handling hard-to-detect cells and improving matching accuracy. Additionally omitting track interpolation with the Kalman Filter further degraded performance, with DET, LNK, and TRA scores at $0.9083$, $0.8570$, and $0.9016$, respectively. By contrast, omitting interpolation alone resulted in only a modest performance decline (DET: $0.9344$, LNK: $0.8982$, TRA: $0.9297$), highlighting that low-confidence detections contribute most strongly to tracking performance.
Next, we analyzed the effect of the maximum number of frames $n$ for which we retain lost cells in the memory bank (see Fig. 3). Retaining cells for at least $5$ frames improved performance from a TRA score of $0.929$ to $0.936$, but extending the memory beyond $5$ frames yielded no further benefit. Notably, keeping cells beyond $5$ frames even led to a decline in TRA performance, likely because outdated candidates complicated the matching process. We observe the same effect of $n$ on DET. For the LNK metric, retaining cells for $5$ frames led to an initial improvement from $0.8982$ to $0.9154$, but extending the memory further had only a limited effect.
@@ 5 Analysis
We next applied TransientTrack to the entire videos, including unannotated frames, to enable a large-scale analysis of cell division behavior and inheritability. Specifically, we processed $154$ videos in total, $36$ each for the control-, low-, and medium-dose groups, and $46$ for the high-dose group. Across these datasets, we tracked 28,60628{,}606 individual cells, identifying 8,6518{,}651 cell divisions and 2,3502{,}350 cell deaths.
@@ 5.1 Treatment Insights from Cell Tracking
In summary, these results illustrate the utility of our approach in quantifying cell dynamics and their response to treatment. By exposing the U2-OS cell line to different doses of cisplatin, we unveiled the overall cell population behavior in response to a standard chemotherapeutic drug. Specifically, the method reveals, a dose-dependent suppression of mitosis, a delayed but robust induction of apoptosis, and the resulting turnover in net population growth.
@@ 6 Discussion & Conclusion
Beyond evaluating tracking performance, we applied TransientTrack in a large-scale lineage analysis of cisplatin-treated cancer cells. In $154$ videos, we reconstructed trajectories of 29,00029{,}000 cells, detected nearly 8,7008{,}700 mitotic events and 2,4002{,}400 apoptotic events. This large-scale study highlighted well-known biological responses: inhibition of mitosis, dose-dependent induction of apoptosis, and the resulting turnover in net population growth. Furthermore, we showed that cell size is not reliably inherited across generations, while sister-cell size correlations persist but decline under treatment. Additionally, lineage progression is strongly curtailed at higher drug concentrations.
@@ 7 Acknowledgments
Florian Bürger and Katarzyna Bozek were supported by the North Rhine-Westphalia return program (311-8.03.03.02-147635), Martim Gomes Dias was supported by the BMBF program Junior Group Consortia in Systems Medicine (01ZX1917B). Noémie Moreau was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project No. 386793560. Florian Bürger, Katarzyna Bozek, Martim Gomes Dias, and Noémie Moreau were hosted by the Center for Molecular Medicine Cologne. We thank the IT Center of the University of Cologne (ITCC) for providing support and computing time.

### v2

@@ 1 Introduction
To evaluate our approach we used a particularly challenging dataset of $309$ time-lapse videos of human osteosarcoma-derived cancer cell-line (U-2 OS) in culture. Each video is $93$ hours long and the cells are tagged with three fluctuating signals, corresponding to the circadian rhythm, cell fitness, and cell cycle. In $117$ videos, cisplatin, a chemotherapeutic drug, was added to induce DNA damage. For a subset of these videos, we manually annotated cell positions, key cellular events (mitosis and apoptosis), and lineage relationships by linking each cell to its parent and daughter cells. After establishing the tracking performance of our method on a subset of annotated videos, we applied it to the full dataset to conduct a large-scale analysis revealing dose-dependent reductions in proliferation, delayed apoptosis onset, and characteristic lineage signatures reflecting how treatment disrupts cell-state inheritance. Together, these results demonstrate how our approach can uncover both global and lineage-specific dynamics in complex imaging datasets.
@@ 3.1 Data
Our dataset examines the proliferation status and cell cycle phase response of single cells to a chemotherapeutic drug cisplatin (granada2020effects). The dataset comprises $309$ videos, each containing $234$ frames at a size of $1024\times 1024$ pixels per frame. Frames were captured at $30$-minute intervals over a total duration of $117$ hours per video using a fluorescence time-lapse microscopy setup. To monitor cellular dynamics, three distinct fluorescent markers — cell cycle reporter hGeminin-CFP, circadian clock reporter NR1D1::VNP, and cell fitness reporter p53-mKate — were employed, resulting in multi-channel images that capture transient, oscillatory signals over time, providing insight into the state and activity of cells. In $118$ of these videos, cisplatin was added after $48$ hours (frame $96$) in three dosage groups: high ($46$ videos, $13$ µM), medium ($36$ videos, $10$ µM), and low ($36$ videos, $7$ µM). Cisplatin affects DNA synthesis and DNA-damage repair, leading to cell division inhibition and ultimately cell death (yu2016cisplatin). We also have $36$ control videos where cisplatin was not added.
@@
We utilized $10$ of the annotated videos to evaluate tracking performance. The remaining annotated videos were used to train the detector. They were randomly split into training, validation, and test sets with a 0.80.8/0.10.1/$0.1$ ratio, respectively. For analysis, we utilized a subset of the whole dataset for which we have drug dosage information ($154$ videos), including all $234$ frames per video. We published the complete dataset at https://doi.org/10.4126/FRL01-006526608.
@@ 3.2.2 Matching
Our matching is carried out sequentially between two consecutive frames. During the first stage of matching, we match high-confidence detections of frame $t$ and $t+1$ with each other. We base our matching on the distance between the centroids of the detections and the similarity of their embeddings. For the distance, we use the Euclidean distance d(c,c^)d($\hat{c}$) which is the distance between the centroids $c_{\text{cent}}$ and $\hat{c}_{\text{cent}}$ of two cells $c$ and $\hat{c}$, respectively. To enforce embedding consistency during matching, we introduce a similarity threshold $\tau^{\text{sim}}$ applied to the $L^{1}$ distance between cell embeddings $c_{\text{emb}}$ and $\hat{c}_{\text{emb}}$:
@@ 4.2.1 CTC Evaluation
Regarding the CTC metrics (Table 1), TransientTrack achieved a DET score of $0.9395$, the LNK measure of $0.9156$ and a TRA score of $0.9364$. For reference, DET scores reported in the CTC challenge range from $0.876$ to $0.997$, LNK scores from $0.947$ to $1.000$, and TRA scores from $0.802$ to $0.994$ across different datasets and their best-performing methods. Although these values are not directly comparable due to substantial differences in the datasets and their difficulty, it is noteworthy that our method achieved scores within a similar range — except for the LNK score — even though it was evaluated on a considerably more challenging dataset.
Figure 2 presents the performance of TransientTrack across different drug dose groups, evaluated using the TRA metric and their respective number of ground truth tracks. The results indicate that our approach achieved consistently high performance across the high-, medium-, and low-dose groups, which contain an average of $62$, $83$, and $102$ tracks per video, respectively. However, performance slightly declines as cell density increases, particularly in the control group. This group includes the fewest videos, yet exhibits the highest average track count per video ($\sim$300). Despite this, TransientTrack maintains strong tracking accuracy, with average TRA scores of $0.9434$, $0.9387$, and $0.9376$ for the low-, medium-, and high-dose groups, respectively.
@@
CTC Metrics DET $0.9395\pm 0.0098$
LNK $0.9156\pm 0.0179$
TRA $0.9364\pm 0.0104$
MOT Metrics HOTA $0.7670\pm 0.0366$
DetA $0.7775\pm 0.0328$
AssA $0.7579\pm 0.0460$
MOTA $0.7963\pm 0.0421$
MOTP $0.9253\pm 0.0055$
IDF1 $0.8161\pm 0.0412$
@@ 4.2.2 MOT Evaluation
TransientTrack showed consistent MOT performance across all evaluated metrics (Table 1). For overall tracking accuracy, as measured by HOTA, it reached $0.767$, with detection accuracy (DetA) of $0.778$ and association accuracy (AssA) of $0.758$. The MOTA score of $0.796$ indicates that most cells were tracked correctly despite the challenging imaging conditions, although some errors from missed detections, false positives, or identity switches still occurred. In contrast, the MOTP score exceeded $0.925$, highlighting that once cells were detected, their positions were estimated with very high spatial precision. The IDF1 score of $0.816$ confirmed that object identities were maintained reliably in time, supporting the construction of accurate lineage trees and the detection of mitotic and apoptotic events. Taken together, these results suggest that TransientTrack excels at precise cell localization while occasionally facing difficulties in maintaining perfect temporal associations, an expected outcome in videos with transient fluorescent markers, heterogeneous signal quality, and frequent apoptotic events.
@@
TransientTrack $\mathbf{0.9395\pm 0.0098}$ $\mathbf{0.9156\pm 0.0179}$ $\mathbf{0.9364\pm 0.0104}$
Only High-Conf. Detections $0.9202\pm 0.0153$ $0.8794\pm 0.0321$ $0.9149\pm 0.0170$
No Kalman Filter $0.9345\pm 0.0132$ $0.8981\pm 0.0266$ $0.9297\pm 0.0144$
None $0.9083\pm 0.0162$ $0.8571\pm 0.0337$ $0.9017\pm 0.0178$
@@ 4.3 Ablations
In order to assess the contribution of various features or our method to the accuracy, we conducted ablation experiments by selectively omitting them (Table 2). We considered four configurations: all features enabled, usage of only high-confidence detections, no track interpolation with Kalman Filter, and neither. In order to evaluate the different settings, we adhered to the CTC metrics only, since they are more suitable for capturing detection, tracking, and mitosis detection. Only using high-confidence detections led to a substantial drop in performance, with the mean scores for the metrics decreasing from $0.9395$ to $0.9202$ (DET), $0.9156$ to $0.8794$ (LNK), and from $0.9364$ to $0.9149$ (TRA). This underscores the importance of low-confidence detections as a fallback mechanism for handling hard-to-detect cells and improving matching accuracy. Additionally omitting track interpolation with the Kalman Filter further degraded performance, with DET, LNK, and TRA scores at $0.9083$, $0.8571$, and $0.9017$, respectively. By contrast, omitting interpolation alone resulted in only a modest performance decline (DET: $0.9345$, LNK: $0.8981$, TRA: $0.9297$), highlighting that low-confidence detections contribute most strongly to tracking performance.
Next, we analyzed the effect of the maximum number of frames $n$ for which we retain lost cells in the memory bank (see Fig. 3). Retaining cells for at least $5$ frames improved performance from a TRA score of $0.9297$ to $0.9364$, but extending the memory beyond $5$ frames yielded no further benefit. Notably, keeping cells beyond $10$ frames even led to a decline in TRA performance, likely because outdated candidates complicated the matching process. We observe the same effect of $n$ on DET. For the LNK metric, retaining cells for $5$ frames led to an initial improvement from $0.8981$ to $0.9156$, but extending the memory further had only a limited effect.
@@ 5 Analysis
We next applied TransientTrack to the entire videos, including unannotated frames, to enable a large-scale analysis of cell division behavior and inheritability. Specifically, we processed $154$ videos in total, $36$ each for the control-, low-, and medium-dose groups, and $46$ for the high-dose group. Across these datasets, we tracked 28,89028{,}890 individual cells, identifying 8,7738{,}773 cell divisions and 2,3452{,}345 cell deaths.
@@ 5.1 Treatment Insights from Cell Tracking
In summary, these results illustrate the utility of our approach in quantifying cell dynamics and their response to treatment. By exposing the U-2 OS cell line to different doses of cisplatin, we unveiled the overall cell population behavior in response to a standard chemotherapeutic drug. Specifically, the method reveals, a dose-dependent suppression of mitosis, a delayed but robust induction of apoptosis, and the resulting turnover in net population growth.
@@ 6 Discussion & Conclusion
Beyond evaluating tracking performance, we applied TransientTrack in a large-scale lineage analysis of cisplatin-treated cancer cells. In $154$ videos, we reconstructed trajectories of 29,00029{,}000 cells, detected nearly 8,8008{,}800 mitotic events and 2,3002{,}300 apoptotic events. This large-scale study highlighted well-known biological responses: inhibition of mitosis, dose-dependent induction of apoptosis, and the resulting turnover in net population growth. Furthermore, we showed that cell size is not reliably inherited across generations, while sister-cell size correlations persist but decline under treatment. Additionally, lineage progression is strongly curtailed at higher drug concentrations.
@@ 7 Acknowledgments
Florian Bürger and Katarzyna Bozek were supported by the North Rhine-Westphalia return program (311-8.03.03.02-147635), Martim Dias Gomes was supported by the BMBF program Junior Group Consortia in Systems Medicine (01ZX1917B). Noémie Moreau was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project No. 386793560. Florian Bürger, Katarzyna Bozek, Martim Dias Gomes, and Noémie Moreau were hosted by the Center for Molecular Medicine Cologne. We thank the IT Center of the University of Cologne (ITCC) for providing support and computing time.